OpenStudy (anonymous):

3^5 / (-3)^7 ... I believe that makes the base 0, right? So then what happens with the exponents? Help please!

jimthompson5910 (jim_thompson5910):

When dividing exponential expressions with the same base, you subtract the corresponding exponents.

jimthompson5910 (jim_thompson5910):

So you first need to make sure that the bases are the same

jimthompson5910 (jim_thompson5910):

Notice how (-3)^7 = (-1*3)^7 (-3)^7 = (-1)^7*(3)^7 (-3)^7 = -(3)^7

OpenStudy (anonymous):

But they aren't, so now I am confused what I do instead. This is just one part of the question

OpenStudy (anonymous):

I need someone to step by step explain what they are doing when they move all these numbers around and such, because I am not undetstanding by just looking at it

jimthompson5910 (jim_thompson5910):

So 3^5/(-3)^7 is the same as 3^5/( -(3)^7 ) which simplifies to - 3^5/3^7

jimthompson5910 (jim_thompson5910):

Do you see how I'm getting all this?

OpenStudy (anonymous):

I'm not because it's just numbers, I need it explained in english word by word step by step because just analyzing it confuses me even more. I need to verbally know what im doing so i can talk myself through the question

jimthompson5910 (jim_thompson5910):

alright

jimthompson5910 (jim_thompson5910):

In going from (-3)^7 to (-1*3)^7, I'm factoring -3 into -1 times 3

jimthompson5910 (jim_thompson5910):

So (-3)^7 = (-1*3)^7

jimthompson5910 (jim_thompson5910):

From there, I'm raising everything inside the parenthesis to the 7th power like so (-3)^7 = (-1)^7*(3)^7

OpenStudy (anonymous):

Why are you doing that, where is that -1 coming from, why did you move the -3 to positive exponent?

jimthompson5910 (jim_thompson5910):

Because I want to extract out the 3 (since the first base is a 3)

jimthompson5910 (jim_thompson5910):

but the second is a -3, so we can say -3 = -1 times 3

OpenStudy (anonymous):

so just mentally you thought of that? is there something that tells you to do that?

jimthompson5910 (jim_thompson5910):

mentally I did (just because I've been doing this a while), but you can divide the first base (3) by the second base (-3) to get 3/(-3) = -1 So you need to factor out a negative one from the second number to get the first number

OpenStudy (anonymous):

Let me explain the full question. It is long, and I'm just focussing ONE of four parts so doing this will take all day it seens? i am not understanding

jimthompson5910 (jim_thompson5910):

alright

OpenStudy (anonymous):

okay so always divide first base by second base? and use that number always? if there are uneven anyway?

OpenStudy (anonymous):

not counting if they are the same as that is a much simpler formula

jimthompson5910 (jim_thompson5910):

well in this case, it works out evenly...but it won't always give you an integer

OpenStudy (anonymous):

i wish i could actually speak to someone on the phone or something, typing is so hard to explain everything

jimthompson5910 (jim_thompson5910):

i understand, but you're doing great...you just have to keep at it

jimthompson5910 (jim_thompson5910):

you can take a screenshot of it to save typing

OpenStudy (anonymous):

i am taking an online class and this is honestly the hardest thing of my life. hardest thing i have EVER done or lack there of i guess you ccan say since i cant do anything :/

jimthompson5910 (jim_thompson5910):

I'm sure you're doing fine. You just have to take things one step at a time.

jimthompson5910 (jim_thompson5910):

and don't be so hard on yourself, I'm sure you can do lots of things well

OpenStudy (anonymous):

It is not letting me attach the picture. can we try to email? can i do that on this site?

jimthompson5910 (jim_thompson5910):

hmm maybe you can post the pic on a 3rd party site (like photobucket) and link it in

OpenStudy (anonymous):

i am trying

OpenStudy (anonymous):

as i said, it is just one part of the equation but i want to focus on just it so i dont get any more confused than i already am

jimthompson5910 (jim_thompson5910):

ok I see it now, one sec

jimthompson5910 (jim_thompson5910):

you agree that you can factor -3 into -1 times 3 right?

OpenStudy (anonymous):

yes that makes sense

jimthompson5910 (jim_thompson5910):

so why not factor -3 into -1 times 3 so in step 2, rewrite everything in step 1, but instead of writing -3, write -1 times 3

OpenStudy (anonymous):

But I did step two, the second line. is that wrong?

jimthompson5910 (jim_thompson5910):

then raise every factor to their corresponding exponent

jimthompson5910 (jim_thompson5910):

no, it's not wrong

OpenStudy (anonymous):

The main question is the top top line. secondf line is what i did

jimthompson5910 (jim_thompson5910):

I'm just breaking it down further

OpenStudy (anonymous):

I dont understand how you want me to do that with everything else around it. what am i doing?

jimthompson5910 (jim_thompson5910):

let me draw it out, one second

jimthompson5910 (jim_thompson5910):

We have the following problem \[\Large \frac{\left( 3x^4y^5z^7\right )^{5}}{\left(-3x^3yz^4 \right )^{7}}\] correct?

OpenStudy (anonymous):

yes, that is the original question and i broke it down further and did one other step, the second line on the picture i sent you

OpenStudy (anonymous):

and frokm what i understand, that was right. now you are talking to me about the third step and that is where i am getting vconfused

jimthompson5910 (jim_thompson5910):

here's what I mean in step 2

jimthompson5910 (jim_thompson5910):

\[\Large \frac{\left( 3x^4y^5z^7\right )^{5}}{\left(-3x^3yz^4 \right )^{7}}\] \[\Large \frac{\left( 3x^4y^5z^7\right )^{5}}{\left(-1*3x^3yz^4 \right )^{7}}\]

jimthompson5910 (jim_thompson5910):

step 3 is then...

OpenStudy (anonymous):

so what you are saying isw that before i do anything else, i should try to get the bases the same, then go forth, right?

jimthompson5910 (jim_thompson5910):

\[\Large \frac{\left( 3x^4y^5z^7\right )^{5}}{\left(-3x^3yz^4 \right )^{7}}\] \[\Large \frac{\left( 3x^4y^5z^7\right )^{5}}{\left(-1*3x^3yz^4 \right )^{7}}\] \[\Large \frac{\left( 3\right)^{5}\left(x^4\right)^{5}\left(y^5\right)^{5}\left(z^7\right )^{5}}{\left(-1\right )^{7}*\left(3\right )^{7}\left(x^3\right )^{7}\left(y\right )^{7}\left(z^4 \right )^{7}}\]

jimthompson5910 (jim_thompson5910):

Do you see where to go from here?

OpenStudy (anonymous):

So let me try to do this but it will be wrong i am assuming so can you help explain why it is wrong?

jimthompson5910 (jim_thompson5910):

alright, go for it

OpenStudy (anonymous):

let me ask you something first

jimthompson5910 (jim_thompson5910):

whats that

OpenStudy (anonymous):

should i just be solving each line as it's own not even acting as if the other exists right niow? until i break it down firther?

OpenStudy (anonymous):

like 3 to the power of five is 243. is that what i should be doing?

jimthompson5910 (jim_thompson5910):

that's a good way to do it, taking it one step at a time (and not focusing on the other steps) will simplify things greatly

jimthompson5910 (jim_thompson5910):

you can have that off to the side

jimthompson5910 (jim_thompson5910):

3^5 = 243 (-1)^7 = -1 etc etc

jimthompson5910 (jim_thompson5910):

once you have all the pieces, you can put them together to form a complete step

OpenStudy (anonymous):

i should leave that for now and simplify the other ones with the exponents in brackets and exponents out of brackets, right? before i do anything else

jimthompson5910 (jim_thompson5910):

this will avoid generating a million steps

jimthompson5910 (jim_thompson5910):

you can think of \[\Large \left( 3\right)^{5}\] as \[\Large \left( 3^{1}\right)^{5}\]

jimthompson5910 (jim_thompson5910):

since 3^1 = 3

jimthompson5910 (jim_thompson5910):

so you can then multiply each outer exponent with the corresponding inner exponent (for each term)

OpenStudy (anonymous):

k but what im sayign is that one is already simplifed to the next level i guess you could call it so i should make the other ones with just one exponent like it is, right? sorry i make no sense. i am the dumbest person ever

jimthompson5910 (jim_thompson5910):

no you're not, don't even think that

OpenStudy (anonymous):

yes that is what i am saying. i think. okay hang on

jimthompson5910 (jim_thompson5910):

yes go ahead and do what you're describing and I'll tell you if you did it correctly or not

OpenStudy (anonymous):

okay i have a feeling the bottom line is wrong but i did my best.

jimthompson5910 (jim_thompson5910):

show me what you got

jimthompson5910 (jim_thompson5910):

alright, let me have a look

jimthompson5910 (jim_thompson5910):

how did you get 3^42 ?

OpenStudy (anonymous):

the 42 is wrong ... 49 instead? forgot how to count lol

OpenStudy (anonymous):

or maybe just leave at 3^7?

jimthompson5910 (jim_thompson5910):

Yes just 3^7 or 2187

jimthompson5910 (jim_thompson5910):

Since 3^7 = 2187

OpenStudy (anonymous):

okay now can you explain why so i remember? because it is confusing me

jimthompson5910 (jim_thompson5910):

\[\Large (3)^7 = \left(3^1\right)^7\] agreed?

OpenStudy (anonymous):

okay see i feel we should keep it at the first step because i didnt know that wheni wads doing the rest of the equation

jimthompson5910 (jim_thompson5910):

you mean keep it (3)^7 ?

OpenStudy (anonymous):

well like it is not being timsed by sseven like all the rest, so i need to remember that or i will always screw it up you kinow?

jimthompson5910 (jim_thompson5910):

ah i see what you mean

jimthompson5910 (jim_thompson5910):

you just have to remember that there's a 1 up there (by default)

OpenStudy (anonymous):

i cant explain it, thats not what i mean. i dont know how to say it

jimthompson5910 (jim_thompson5910):

if there is no exponent

jimthompson5910 (jim_thompson5910):

hmm alright, but just keep that in mind

OpenStudy (anonymous):

i dont understand what you said. even more confusion. you dont have to help mne if this is too much for you

jimthompson5910 (jim_thompson5910):

I'm sorry, I didn't mean to confuse you further

OpenStudy (anonymous):

I just want to keep that 1 there so i know, because if not i am doing something i should not be doing it seems

jimthompson5910 (jim_thompson5910):

alright

OpenStudy (anonymous):

anyway, now that i have that step, with the number changed from 42 to a simple 7, i am trying to figure out what to do next

jimthompson5910 (jim_thompson5910):

your next step is to divide and simplify

OpenStudy (anonymous):

how do i divide like this? i dont really divide, i add right? or no, subtract actually